OpenStudy (anonymous):

64y^2-x^3y^2=(64-x^3)

OpenStudy (campbell_st):

well if you are being asked to solve it... factor the left hand side... the common factor is y^2... so what would it look like if you factored...?

OpenStudy (anonymous):

y^2(64-x^3) y^2(4^3-x^3) The rest ?

OpenStudy (campbell_st):

yep... so your equation is y^2(64 - x^3) = (64 - x^3) so what do you do next....?

OpenStudy (anonymous):

No clue what to do next

OpenStudy (campbell_st):

so are you being asked to solve the equation... or something else..?

OpenStudy (anonymous):

Its asking me 64y^4-x^3y^4= ? then it asks me if the polynomial is prime or not?

OpenStudy (campbell_st):

well it looks like they have divided by y^4 and left 64 - x^3 then, based in the information, its saying... can it be factored further... if it can't be factored its prime but 64 -x^3 can be factored to (4 - x)(16 + 4x + x^2)

OpenStudy (campbell_st):

the question doesn't make a lot of sense to me... sorry

OpenStudy (anonymous):

it's #7

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well the solution you have on the paper is correct y^3(4 -x)(16 + 4x + x^2)

OpenStudy (campbell_st):

64 - x^3 is the difference of 2 cubes as it becomes 4^3 - x^3 the standard form is a^3 - b^3 = (a -b)(a^2 +ab + b^2)